OpenStudy (anonymous):

Evaluate the expression below when x = 5. 3x(2x - 7) Select one of the options below as your answer: A. 0 B. 24 C. 45 D. 143

OpenStudy (anonymous):

The X Is A Variable

OpenStudy (anonymous):

C

OpenStudy (anonymous):

3(5)(2(5)-7)=15(10-7)=15(3)=45 (C)